# Detection of Gravitational Anomaly at Low Acceleration from a Highest-quality Sample of 36 Wide Binaries with Accurate 3D Velocities

### v1

@@ IV.3 Exploring possible systematic errors: Can the gravitational anomaly be removed?
The Gaia limiting magnitude of 20.5 out to the limit of our sample of 150 pc implies that stellar companions at separations beyond the Gaia resolution of 0.3 arcsec can be excluded as contaminating our sample. At closer separations, the combination of only 5% close companions identified by our Speckle campaign and the clearing of a subsample of our final clean sample through this technique, exclude a further range of potential stellar companions. In order to bring our results into consistency with Newtonian expectations we would require, as detailed above, about 9 of our binaries to be contaminated, something which would require a probability of about $(0.05)^{9}$ if unresolved stellar companions were to blame. At even closer separation, where close binary periods become comparable to the Gaia DR3 temporal range of 34 months, the use of ruwe restrictions for all (e.g. see Castro-Ginard et al. (2024) where the presence of unresolved stellar companions of all types has been shown to be incompatible with stars in DR3 for values of $\texttt{ruwe}>1.15$, a limit only one of our binaries in the clean sample crosses) and consistent multi-epoch high quality radial velocity measurements, in practice eliminate stellar companion contamination as a concern.
@@ VI Summary, Conclusion, and outlook
We started from assembling a new collection of 306 Gaia DR3 wide binaries with relatively precise radial velocities. Speckle interferometric imaging of 414 wide binaries of similar qualities shows that about 5% can have resolvable faint companions. From the initial collection, we carefully selected a high-quality sample of 75 isolated wide binaries, based on limits such as distance <150<150 pc from the Sun, Gaia’s $\texttt{ruwe}<1.25$, a narrow locus in the CM diagram, and relative radial velocity ($v_{r}$) error <100<100 m s-1.
@@ E.2 Data Processing
The first step of the data processing is the dark field correction of detected images $i^{\prime}_{n}(\vec{x})$:
@@
where $i_{n}(\vec{x})$ is the intensity at point $(\vec{x})$ in the corrected image, $Dark(\vec{x})$ is the average dark image captured with closed shutter. Then for each specklegram we calculate the following criteria:
@@
where the summations extend over the complete n-th specklegram (McCarthy and Cobb, 1986). Speckle images are then ranked by their Sh1, where the ”best” frames are those with the lowest Sh1. Two images in Figure 22 show the best (a) and worst (b) speckle images in a series of 2000 specklegrams.
@@ E.3 High-Resolution
Then, applying the Fourier transform to equation E3
@@
where $I_{n}(\vec{\omega})$ and $O(\vec{\omega})$ are the Fourier transform of the image and object intensity distributions respectively, and $OTF_{n}(\vec{\omega})$ is the Optical Transfer Function (OTF), which is the equivalent of the $PSF$ in Fourier space. The $OTF_{n}(\vec{\omega})$ is the same for all objects within the isoplanatic region. Equation E4 is the base for all high-resolution techniques. We applied various image processing techniques to reconstruct the power spectrum and obtain a high-resolution image of the star. This involves analyzing the speckle patterns in each short-exposure frame and combining them.
@@ E.4 Power Spectrum
In case of low light images, the averaged power spectrum can be expressed as (Kerp et al., 1992):
@@
where $P(\vec{\omega})$ is the unshifted estimation of the power spectrum, $q$ is some constant, $\left|G(\vec{\omega})\right|^{2}$ is the power spectrum of the photon event shape function also known as photon bias. The photon bias $\left|G(\vec{\omega})\right|^{2}$ can be determined as the normalized power spectrum of the night sky. The $\left|G(\vec{\omega})\right|^{2}$ is constant in Y direction for this camera. Thus, it can be determined directly from $PS(\vec{\omega})$ (Figure 23, (a)) by analysis of it’s part beyond the cut-off frequency of telescope. The unshifted power spectrum of specklegrams $P({\bf f})$ shown in Figure 23 (b). Therefore, it can be presented as:
@@
[…]
where $\left|O(\vec{\omega})\right|^{2}$ power spectrum of the object, $\left\langle\left|S_{n}(\vec{\omega})\right|^{2}\right\rangle$ is the speckle interferometric transfer function. The speckle interferometric transfer function can be obtained by observing of a reference star or one can construct universal synthetic speckle interferometric transfer function (Tokovinin et al., 2010). Since we are concerned only with the presence or absence of a close stellar component, this information can be obtained directly from $P(\vec{\omega})$ without requiring the speckle interferometric transfer function. In particular, we calculated the high resolution autocorrelation function in polar coordinates $ACF_{p}$:
@@
where $W(r,\phi)$ is the window which excludes part of $P(r,\phi)$ beyond the cut-off frequency of telescope $f_{T}$ and for frequencies lower atmospheric cutoff $f_{A}$. Also taking in to account central symmetry of $P(r,\phi)$ equation E8 can be rewritten as:
@@
A star has a component if the high-resolution $ACF_{p}$ has a pronounced maximum. $ACF_{p}$ allows finding the component even when the power spectrum is distorted by vibrations and strong aberrations of the telescope (Orlov, 2021). One example of $ACF_{p}$ is shown in figure 23 (c).
@@ E.5 Speckle Holography
Figure 24: The typical wide binary system, composed of Gaia DR3 1172915990414659328 (G = 9.2) and Gaia DR3 1172920487244742912 (G = 9.4). (a) High-resolution autocorrelation of Gaia DR3 1172915990414659328. (b) High-resolution speckle holography image of Gaia DR3 1172915990414659328. (c) High-resolution autocorrelation of Gaia DR3 1172920487244742912. (d) High-resolution speckle holography image of Gaia DR3 1172920487244742912.
For most wide binaries we found no hidden stellar components (Figure 24). However, during data processing we identified several interesting systems — for example, Gaia DR3 4493544185516358656, which turned out to be a triple star system (Fig. 25). To investigate such cases in more detail, we applied speckle holography to the entire dataset. The method can be summarized as follows:
If one has a point source $O^{r}(\vec{x})=\delta(\vec{x}=\vec{x^{r}})$ inside this area, the equation E4 for this star becomes:
@@
Substitution of equation E10 into equation E4 gives:
@@
Figure 25: The triple system ABC, composed of Gaia DR3 4493544185516358656 (G = 12.86), Gaia DR3 4493544082437142400 (G = 13.26), and Gaia DR3 4493544082437142784 (G = 14.13). (a) Power spectrum of subsystem BC. (b) High-resolution autocorrelation of subsystem BC. (c) High-resolution speckle holography image. It was derived from Speckle Interferometry (Labeyrie, 1970), being the main difference that speckle holography works to remove atmospheric aberrations from each short exposure image and the phase information is kept. The separation between components B and C is 0.1″(8 pc).
@@
At large separations from the primary star (more than about three times the seeing disk) the noise reaches a steady level, and the detection limit becomes roughly constant. At smaller separations, however, it strongly depends on angular distance from the primary star. This behavior is summarized with a contrast curve (Figure 26). The contrast curve plots the faintest detectable companion (in terms of brightness difference, $\Delta m$) as a function of separation from the primary star. In other words, it tells you: If there had been a companion this bright at this distance, we would have seen it.
Figure 26: The $5\sigma$ Contrast curves. (a) The curve is for a star with a magnitude of 12, while (b) is the curve for a star with a magnitude of 10.
Since no companions were detected in the vast majority of cases, we present for each star only two representative points of the contrast curve, corresponding $\Delta m=3$ and $\Delta m=4$.
In total, we obtained 1019 measurements for 391 selected wide binaries. Many targets were observed two or more times, allowing us to retain only the highest-quality measurements for each system. The table 10 lists the results for 782 stars.

### v2

@@ IV.3 Exploring possible systematic errors: Can the gravitational anomaly be removed?
The Gaia limiting magnitude of 20.5 out to the limit of our sample of 150 pc implies that stellar companions at separations beyond the Gaia resolution of $\approx 0.5$ arcsec (Gaia Collaboration et al., 2018; Lindegren et al., 2021) can be excluded as contaminating our sample. At closer separations, the combination of only 5% close companions identified by our Speckle campaign and the clearing of a subsample of our final clean sample through this technique, exclude a further range of potential stellar companions. In order to bring our results into consistency with Newtonian expectations we would require, as detailed above, about 9 of our binaries to be contaminated, something which would require a probability of about $(0.05)^{9}$ if unresolved stellar companions were to blame. At even closer separation, where close binary periods become comparable to the Gaia DR3 temporal range of 34 months, the use of ruwe restrictions for all (e.g. see Castro-Ginard et al. (2024) where the presence of unresolved stellar companions of all types has been shown to be incompatible with stars in DR3 for values of $\texttt{ruwe}>1.15$, a limit only one of our binaries in the clean sample crosses) and consistent multi-epoch high quality radial velocity measurements, in practice eliminate stellar companion contamination as a concern.
@@ VI Summary, Conclusion, and outlook
We started from assembling a new collection of 306 Gaia DR3 wide binaries with relatively precise radial velocities. Speckle interferometric imaging of 391 wide binaries of similar qualities shows that about 5% can have resolvable faint companions. From the initial collection, we carefully selected a high-quality sample of 75 isolated wide binaries, based on limits such as distance <150<150 pc from the Sun, Gaia’s $\texttt{ruwe}<1.25$, a narrow locus in the CM diagram, and relative radial velocity ($v_{r}$) error <100<100 m s-1.
@@ E.2 Data Processing
The first step of data processing is the dark field correction of detected images $i^{\prime}_{n}(\vec{x})$:
@@
where $i_{n}(\vec{x})$ is the intensity at point $\vec{x}$ in the corrected image, $Dark(\vec{x})$ is the average dark image captured with the closed shutter. Then for each specklegram we calculate the following criteria:
@@
where the summations extend over the complete n-th specklegram (McCarthy and Cobb, 1986). Speckle images are then ranked by their $Sh1$, where the “best” frames are those with the lowest $Sh1$. Two images in Figure 22 show the best (a) and worst (b) speckle images in a series of 2000 specklegrams.
@@ E.3 High-Resolution
Then, applying the Fourier transform to Equation (E3)
@@
where $I_{n}(\vec{\omega})$ and $O(\vec{\omega})$ are the Fourier transform of the image and object intensity distributions respectively, and $OTF_{n}(\vec{\omega})$ is the Optical Transfer Function (OTF), which is the equivalent of the $PSF$ in Fourier space. The $OTF_{n}(\vec{\omega})$ is the same for all objects within the isoplanatic region. Equation (E4) is the base for all high-resolution techniques. We applied various image processing techniques to reconstruct the power spectrum and obtain a high-resolution image of the star. This involves analyzing the speckle patterns in each short-exposure frame and combining them.
@@ E.4 Power Spectrum
[…]
In the case of low light images, the averaged power spectrum can be expressed as (Kerp et al., 1992):
@@
where $P(\vec{\omega})$ is the unshifted estimation of the power spectrum, $q$ is some constant, and $\left|G(\vec{\omega})\right|^{2}$ is the power spectrum of the photon event shape function also known as photon bias. The photon bias $\left|G(\vec{\omega})\right|^{2}$ can be determined as the normalized power spectrum of the night sky. The power spectrum $\left|G(\vec{\omega})\right|^{2}$ is constant in the vertical direction for this camera. Thus, it can be determined directly from $PS(\vec{\omega})$ (Figure 23(a)) by analyzing its part beyond the cut-off frequency of the telescope. The unshifted power spectrum of specklegrams $P({\bf f})$ is shown in Figure 23(b). Therefore, it can be presented as:
@@
where $\left|O(\vec{\omega})\right|^{2}$ is the power spectrum of the object, and $\left\langle\left|S_{n}(\vec{\omega})\right|^{2}\right\rangle$ is the speckle interferometric transfer function. The speckle interferometric transfer function can be obtained by observing a reference star, or one can construct universal synthetic speckle interferometric transfer function (Tokovinin et al., 2010).
We are concerned only with the presence or absence of a close stellar component, and the information can be obtained directly from $P(\vec{\omega})$ without requiring the speckle interferometric transfer function. In particular, we calculated the high resolution autocorrelation function in polar coordinates $ACF_{p}$:
@@
where $W(r,\phi)$ is the window that excludes part of $P(r,\phi)$ beyond the cut-off frequency of the telescope $f_{T}$ and for frequencies lower than the atmospheric cutoff $f_{A}$. Also taking into account central symmetry of $P(r,\phi)$, Equation  (E8) can be rewritten as:
@@
A star has a component if the high-resolution $ACF_{p}$ has a pronounced maximum. $ACF_{p}$ allows one to find the component even when the power spectrum is distorted by vibrations and strong aberrations of the telescope (Orlov, 2021). An example of $ACF_{p}$ is shown in Figure 23(c).
@@ E.5 Speckle Holography
Figure 24: The typical wide binary system, composed of Gaia DR3 1172915990414659328 (G = 9.2) and Gaia DR3 1172920487244742912 (G = 9.4). (a) High-resolution autocorrelation of Gaia DR3 1172915990414659328. (b) High-resolution speckle holography image of Gaia DR3 1172915990414659328. (c) High-resolution autocorrelation of Gaia DR3 1172920487244742912. (d) High-resolution speckle holography image of Gaia DR3 1172920487244742912. Panels (a), (b), (c) and (d) are from left to right.
For most wide binaries, hidden stellar components were not detected (Figure 24). However, during data processing we identified several interesting systems — for example, Gaia DR3 4493544185516358656, which turned out to be a triple star system (Figure 25). To investigate such cases in more detail, we applied speckle holography to the entire dataset. The method can be summarized as follows.
If one has a point source $O^{r}(\vec{x})=\delta(\vec{x}=\vec{x^{r}})$ inside the area, Equation (E4) for this star becomes:
@@
Substitution of Equation (E10) into Equation (E4) gives:
@@
Figure 25: The triple system ABC, composed of Gaia DR3 4493544185516358656 (G = 12.86), Gaia DR3 4493544082437142400 (G = 13.26), and Gaia DR3 4493544082437142784 (G = 14.13). (a) Power spectrum of subsystem BC. (b) High-resolution autocorrelation of subsystem BC. (c) High-resolution speckle holography image. It was derived from Speckle Interferometry (Labeyrie, 1970), the main difference being that speckle holography works so as to remove atmospheric aberrations from each short exposure image and the phase information is kept. The separation between components B and C is 0.1″(8 au). Panels (a), (b) and (c) are from left to right.
@@
At large separations from the primary star (more than about three times the seeing disk) the noise reaches a steady level, and the detection limit becomes roughly constant. At smaller separations, however, it strongly depends on angular distance from the primary star. This behavior is summarized with a contrast curve (Figure 26). The contrast curve plots the faintest detectable companion (in terms of brightness difference, $\Delta m$) as a function of separation from the primary star. In other words, it tells: If there had been a companion this bright at this distance, we would have seen it.
Figure 26: The $5\sigma$ Contrast curves. The contrast curve for Gaia DR3 1331462302965590144 (”Too weak”) with a magnitude of 14.3 (left) and the curve for Gaia DR3 4197023904014523776 with a magnitude of 8.57 (right).
Since no companions were detected in the vast majority of cases, for each star we present only two representative points of the contrast curve, corresponding $\Delta m=3$ and $\Delta m=4$.
In total, we obtained 1019 measurements for 391 selected wide binaries. Many targets were observed two or more times, allowing us to retain only the highest-quality measurements for each system. Table 10 lists the results for 782 stars.